# Sequential Conformal Risk Control UQ Uncertainty Quantification

### v1

@@ Appendix F Algorithms
mix box_count_threshold thresholded lac additive 26.3792 0.0188 4.5749 0.0460 17.3165 0.0458 0.0857
mix box_count_threshold thresholded lac multiplicative 26.3792 0.0188 2.5006 0.0516 17.3165 0.0458 0.0946
mix box_count_threshold thresholded aps additive 26.3792 0.0188 4.5749 0.0440 23.4177 0.0415 0.0788
mix box_count_threshold thresholded aps multiplicative 26.3792 0.0188 2.1484 0.0484 16.2460 0.0491 0.0939
mix box_count_threshold pixelwise lac additive 26.3792 0.0188 1.5091 0.0484 13.4423 0.0476 0.0850
mix box_count_threshold pixelwise lac multiplicative 26.3792 0.0188 1.0047 0.0510 13.4423 0.0476 0.0942
mix box_count_threshold pixelwise aps additive 26.3792 0.0188 1.5091 0.0484 14.3149 0.0484 0.0866
mix box_count_threshold pixelwise aps multiplicative 26.3792 0.0188 1.0024 0.0510 12.3572 0.0484 0.0933
mix box_count_threshold boxwise lac additive 26.3792 0.0188 1.8151 0.0484 12.4177 0.0484 0.0860
mix box_count_threshold boxwise lac multiplicative 26.3792 0.0188 1.2427 0.0510 12.4177 0.0484 0.0942
mix box_count_threshold boxwise aps additive 26.3792 0.0188 1.8151 0.0484 13.7468 0.0484 0.0876
mix box_count_threshold boxwise aps multiplicative 26.3792 0.0188 1.2427 0.0510 12.4843 0.0484 0.0942
mix box_count_recall thresholded lac additive 18.8680 0.0186 4.5749 0.0460 17.3165 0.0458 0.0857
mix box_count_recall thresholded lac multiplicative 18.8680 0.0186 2.5006 0.0516 17.3165 0.0458 0.0946
mix box_count_recall thresholded aps additive 18.8680 0.0186 4.5749 0.0440 23.4177 0.0415 0.0788
mix box_count_recall thresholded aps multiplicative 18.8680 0.0186 2.1484 0.0484 16.2460 0.0491 0.0939
mix box_count_recall pixelwise lac additive 18.8680 0.0186 1.5091 0.0484 13.4423 0.0476 0.0850
mix box_count_recall pixelwise lac multiplicative 18.8680 0.0186 1.0047 0.0510 13.4423 0.0476 0.0942
mix box_count_recall pixelwise aps additive 18.8680 0.0186 1.5091 0.0484 14.3149 0.0484 0.0866
mix box_count_recall pixelwise aps multiplicative 18.8680 0.0186 1.0024 0.0510 12.3572 0.0484 0.0933
mix box_count_recall boxwise lac additive 18.8680 0.0186 1.8151 0.0484 12.4177 0.0484 0.0860
mix box_count_recall boxwise lac multiplicative 18.8680 0.0186 1.2427 0.0510 12.4177 0.0484 0.0942
mix box_count_recall boxwise aps additive 18.8680 0.0186 1.8151 0.0484 13.7468 0.0484 0.0876
mix box_count_recall boxwise aps multiplicative 18.8680 0.0186 1.2427 0.0510 12.4843 0.0484 0.0942
giou box_count_threshold thresholded lac additive 26.3792 0.0188 35.0661 0.0451 47.6759 0.0486 0.0879
giou box_count_threshold thresholded lac multiplicative 26.3792 0.0188 29.1481 0.0502 47.6759 0.0486 0.0940
giou box_count_threshold thresholded aps additive 26.3792 0.0188 35.0661 0.0443 46.7573 0.0460 0.0840
giou box_count_threshold thresholded aps multiplicative 26.3792 0.0188 28.3221 0.0490 45.7192 0.0460 0.0903
giou box_count_threshold pixelwise lac additive 26.3792 0.0188 30.3110 0.0476 47.9874 0.0484 0.0888
giou box_count_threshold pixelwise lac multiplicative 26.3792 0.0188 27.6434 0.0499 47.9874 0.0484 0.0930
giou box_count_threshold pixelwise aps additive 26.3792 0.0188 30.3110 0.0476 46.7573 0.0460 0.0862
giou box_count_threshold pixelwise aps multiplicative 26.3792 0.0188 27.4804 0.0490 45.7192 0.0460 0.0903
giou box_count_threshold boxwise lac additive 26.3792 0.0188 36.2020 0.0476 46.8872 0.0476 0.0871
giou box_count_threshold boxwise lac multiplicative 26.3792 0.0188 30.1844 0.0499 46.8872 0.0476 0.0923
giou box_count_threshold boxwise aps additive 26.3792 0.0188 36.2020 0.0476 47.0217 0.0460 0.0855
giou box_count_threshold boxwise aps multiplicative 26.3792 0.0188 30.0534 0.0490 45.7192 0.0460 0.0903
giou box_count_recall thresholded lac additive 18.8680 0.0186 35.0661 0.0451 47.6759 0.0486 0.0879
giou box_count_recall thresholded lac multiplicative 18.8680 0.0186 29.1481 0.0502 47.6759 0.0486 0.0940
giou box_count_recall thresholded aps additive 18.8680 0.0186 35.0661 0.0443 46.7573 0.0460 0.0840
giou box_count_recall thresholded aps multiplicative 18.8680 0.0186 28.3221 0.0490 45.7192 0.0460 0.0903
giou box_count_recall pixelwise lac additive 18.8680 0.0186 30.3110 0.0476 47.9874 0.0484 0.0888
giou box_count_recall pixelwise lac multiplicative 18.8680 0.0186 27.6434 0.0499 47.9874 0.0484 0.0930
giou box_count_recall pixelwise aps additive 18.8680 0.0186 30.3110 0.0476 46.7573 0.0460 0.0862
giou box_count_recall pixelwise aps multiplicative 18.8680 0.0186 27.4804 0.0490 45.7192 0.0460 0.0903
giou box_count_recall boxwise lac additive 18.8680 0.0186 36.2020 0.0476 46.8872 0.0476 0.0871
giou box_count_recall boxwise lac multiplicative 18.8680 0.0186 30.1844 0.0499 46.8872 0.0476 0.0923
giou box_count_recall boxwise aps additive 18.8680 0.0186 36.2020 0.0476 47.0217 0.0460 0.0855
giou box_count_recall boxwise aps multiplicative 18.8680 0.0186 30.0534 0.0490 45.7192 0.0460 0.0903
hausdorff box_count_threshold thresholded lac additive 26.3792 0.0188 1.6257 0.0452 44.6030 0.0503 0.0966
hausdorff box_count_threshold thresholded lac multiplicative 26.3792 0.0188 1.0562 0.0509 44.6030 0.0503 0.1000
hausdorff box_count_threshold thresholded aps additive 26.3792 0.0188 1.6257 0.0444 44.2706 0.0476 0.0849
hausdorff box_count_threshold thresholded aps multiplicative 26.3792 0.0188 1.0562 0.0499 44.2706 0.0476 0.0915
hausdorff box_count_threshold pixelwise lac additive 26.3792 0.0188 1.0936 0.0476 44.6030 0.0503 0.0987
hausdorff box_count_threshold pixelwise lac multiplicative 26.3792 0.0188 1.0006 0.0508 44.6030 0.0503 0.1000
hausdorff box_count_threshold pixelwise aps additive 26.3792 0.0188 1.0936 0.0476 44.2706 0.0476 0.0853
hausdorff box_count_threshold pixelwise aps multiplicative 26.3792 0.0188 1.0006 0.0499 44.2706 0.0476 0.0915
hausdorff box_count_threshold boxwise lac additive 26.3792 0.0188 1.5783 0.0491 44.6030 0.0503 0.0987
hausdorff box_count_threshold boxwise lac multiplicative 26.3792 0.0188 1.0562 0.0509 44.6030 0.0503 0.1000
hausdorff box_count_threshold boxwise aps additive 26.3792 0.0188 1.5783 0.0491 44.2706 0.0476 0.0853
hausdorff box_count_threshold boxwise aps multiplicative 26.3792 0.0188 1.0562 0.0499 44.2706 0.0476 0.0915
hausdorff box_count_recall thresholded lac additive 18.8680 0.0186 1.6257 0.0452 44.6030 0.0503 0.0966
hausdorff box_count_recall thresholded lac multiplicative 18.8680 0.0186 1.0562 0.0509 44.6030 0.0503 0.1000
hausdorff box_count_recall thresholded aps additive 18.8680 0.0186 1.6257 0.0444 44.2706 0.0476 0.0849
hausdorff box_count_recall thresholded aps multiplicative 18.8680 0.0186 1.0562 0.0499 44.2706 0.0476 0.0915
hausdorff box_count_recall pixelwise lac additive 18.8680 0.0186 1.0936 0.0476 44.6030 0.0503 0.0987
hausdorff box_count_recall pixelwise lac multiplicative 18.8680 0.0186 1.0006 0.0508 44.6030 0.0503 0.1000
hausdorff box_count_recall pixelwise aps additive 18.8680 0.0186 1.0936 0.0476 44.2706 0.0476 0.0853
hausdorff box_count_recall pixelwise aps multiplicative 18.8680 0.0186 1.0006 0.0499 44.2706 0.0476 0.0915
hausdorff box_count_recall boxwise lac additive 18.8680 0.0186 1.5783 0.0491 44.6030 0.0503 0.0987
hausdorff box_count_recall boxwise lac multiplicative 18.8680 0.0186 1.0562 0.0509 44.6030 0.0503 0.1000
hausdorff box_count_recall boxwise aps additive 18.8680 0.0186 1.5783 0.0491 44.2706 0.0476 0.0853
hausdorff box_count_recall boxwise aps multiplicative 18.8680 0.0186 1.0562 0.0499 44.2706 0.0476 0.0915
lac box_count_threshold thresholded lac additive 26.3792 0.0188 19.5160 0.0476 1.0651 0.0503 0.0828
lac box_count_threshold thresholded lac multiplicative 26.3792 0.0188 15.2387 0.0499 1.0651 0.0503 0.0909
lac box_count_threshold thresholded aps additive 26.3792 0.0188 19.5160 0.0476 1.0440 0.0484 0.0819
lac box_count_threshold thresholded aps multiplicative 26.3792 0.0188 15.2387 0.0499 1.0440 0.0484 0.0909
lac box_count_threshold pixelwise lac additive 26.3792 0.0188 17.3935 0.0491 1.0651 0.0503 0.0834
lac box_count_threshold pixelwise lac multiplicative 26.3792 0.0188 13.0506 0.0509 1.0651 0.0503 0.0909
lac box_count_threshold pixelwise aps additive 26.3792 0.0188 17.3935 0.0491 1.0440 0.0484 0.0819
lac box_count_threshold pixelwise aps multiplicative 26.3792 0.0188 13.0506 0.0509 1.0440 0.0484 0.0909
lac box_count_threshold boxwise lac additive 26.3792 0.0188 19.5160 0.0476 1.0651 0.0503 0.0828
lac box_count_threshold boxwise lac multiplicative 26.3792 0.0188 15.2387 0.0499 1.0651 0.0503 0.0909
lac box_count_threshold boxwise aps additive 26.3792 0.0188 19.5160 0.0476 1.0440 0.0484 0.0819
lac box_count_threshold boxwise aps multiplicative 26.3792 0.0188 15.2387 0.0499 1.0440 0.0484 0.0909
lac box_count_recall thresholded lac additive 18.8680 0.0186 19.5160 0.0476 1.0651 0.0503 0.0828
lac box_count_recall thresholded lac multiplicative 18.8680 0.0186 15.2387 0.0499 1.0651 0.0503 0.0909
lac box_count_recall thresholded aps additive 18.8680 0.0186 19.5160 0.0476 1.0440 0.0484 0.0819
lac box_count_recall thresholded aps multiplicative 18.8680 0.0186 15.2387 0.0499 1.0440 0.0484 0.0909
lac box_count_recall pixelwise lac additive 18.8680 0.0186 17.3935 0.0491 1.0651 0.0503 0.0834
lac box_count_recall pixelwise lac multiplicative 18.8680 0.0186 13.0506 0.0509 1.0651 0.0503 0.0909
lac box_count_recall pixelwise aps additive 18.8680 0.0186 17.3935 0.0491 1.0440 0.0484 0.0819
lac box_count_recall pixelwise aps multiplicative 18.8680 0.0186 13.0506 0.0509 1.0440 0.0484 0.0909
lac box_count_recall boxwise lac additive 18.8680 0.0186 19.5160 0.0476 1.0651 0.0503 0.0828
lac box_count_recall boxwise lac multiplicative 18.8680 0.0186 15.2387 0.0499 1.0651 0.0503 0.0909
lac box_count_recall boxwise aps additive 18.8680 0.0186 19.5160 0.0476 1.0440 0.0484 0.0819
lac box_count_recall boxwise aps multiplicative 18.8680 0.0186 15.2387 0.0499 1.0440 0.0484 0.0909
TABLE VI: Experimental Results using the DETR-50 model for $\alpha_{tot}=0.2$

### v2

@@ Appendix F Effect of Object Detection and Conformal Settings
mix box_count_threshold thresholded lac additive 25.5884 0.022 4.5646 0.0472 12.7719 0.0449 0.0858
mix box_count_threshold thresholded lac multiplicative 25.5884 0.022 2.717 0.0372 12.7719 0.0449 0.0791
mix box_count_threshold thresholded aps additive 25.5884 0.022 4.5646 0.0416 17.9396 0.037 0.0727
mix box_count_threshold thresholded aps multiplicative 25.5884 0.022 2.2074 0.0404 11.1172 0.0459 0.0818
mix box_count_threshold pixelwise lac additive 25.5884 0.022 1.4788 0.0488 8.228 0.0464 0.083
mix box_count_threshold pixelwise lac multiplicative 25.5884 0.022 1.3337 0.0474 8.228 0.0464 0.0835
mix box_count_threshold pixelwise aps additive 25.5884 0.022 1.4788 0.0488 11.1172 0.0459 0.0829
mix box_count_threshold pixelwise aps multiplicative 25.5884 0.022 1.3337 0.0474 11.1171 0.0459 0.0835
mix box_count_threshold boxwise lac additive 25.5884 0.022 3.8942 0.0464 8.228 0.0464 0.0799
mix box_count_threshold boxwise lac multiplicative 25.5884 0.022 2.086 0.0468 8.228 0.0464 0.0838
mix box_count_threshold boxwise aps additive 25.5884 0.022 3.8942 0.0464 11.1172 0.0459 0.0794
mix box_count_threshold boxwise aps multiplicative 25.5884 0.022 2.086 0.0468 11.1171 0.0459 0.0837
mix box_count_recall thresholded lac additive 17.778 0.0185 4.7833 0.0468 12.0055 0.0471 0.0847
mix box_count_recall thresholded lac multiplicative 17.778 0.0185 2.6168 0.0324 12.0055 0.0471 0.0744
mix box_count_recall thresholded aps additive 17.778 0.0185 4.7833 0.0468 14.9545 0.0468 0.0844
mix box_count_recall thresholded aps multiplicative 17.778 0.0185 2.6168 0.0324 14.9545 0.0468 0.0743
mix box_count_recall pixelwise lac additive 17.778 0.0185 1.7924 0.0475 12.0055 0.0471 0.0807
mix box_count_recall pixelwise lac multiplicative 17.778 0.0185 1.5503 0.0443 12.0055 0.0471 0.0802
mix box_count_recall pixelwise aps additive 17.778 0.0185 1.7924 0.0475 14.9545 0.0468 0.0809
mix box_count_recall pixelwise aps multiplicative 17.778 0.0185 1.5503 0.0443 14.9545 0.0468 0.0805
mix box_count_recall boxwise lac additive 17.778 0.0185 4.1997 0.0476 12.0055 0.0471 0.0798
mix box_count_recall boxwise lac multiplicative 17.778 0.0185 2.4471 0.0432 12.0055 0.0471 0.0801
mix box_count_recall boxwise aps additive 17.778 0.0185 4.1997 0.0476 14.9545 0.0468 0.0799
mix box_count_recall boxwise aps multiplicative 17.778 0.0185 2.4471 0.0432 14.9545 0.0468 0.0804
hausdorff box_count_threshold thresholded lac additive 25.5884 0.022 3.1521 0.0492 42.5629 0.0539 0.1001
hausdorff box_count_threshold thresholded lac multiplicative 25.5884 0.022 1.5522 0.0456 42.5629 0.0539 0.0971
hausdorff box_count_threshold thresholded aps additive 25.5884 0.022 3.1521 0.0492 41.8463 0.0548 0.1009
hausdorff box_count_threshold thresholded aps multiplicative 25.5884 0.022 1.5522 0.0456 41.8461 0.0548 0.0978
hausdorff box_count_threshold pixelwise lac additive 25.5884 0.022 1.047 0.052 42.5629 0.0539 0.099
hausdorff box_count_threshold pixelwise lac multiplicative 25.5884 0.022 1.0425 0.0466 42.5629 0.0539 0.0947
hausdorff box_count_threshold pixelwise aps additive 25.5884 0.022 1.047 0.052 41.8463 0.0548 0.1
hausdorff box_count_threshold pixelwise aps multiplicative 25.5884 0.022 1.0425 0.0466 41.846 0.0548 0.0956
hausdorff box_count_threshold boxwise lac additive 25.5884 0.022 2.6066 0.0506 42.5629 0.0539 0.0968
hausdorff box_count_threshold boxwise lac multiplicative 25.5884 0.022 1.5036 0.049 42.5629 0.0539 0.0971
hausdorff box_count_threshold boxwise aps additive 25.5884 0.022 2.6066 0.0506 41.8463 0.0548 0.0977
hausdorff box_count_threshold boxwise aps multiplicative 25.5884 0.022 1.5036 0.049 41.846 0.0548 0.0981
hausdorff box_count_recall thresholded lac additive 17.778 0.0185 3.8291 0.0472 38.134 0.0515 0.0948
hausdorff box_count_recall thresholded lac multiplicative 17.778 0.0185 1.8169 0.0396 38.134 0.0515 0.0883
hausdorff box_count_recall thresholded aps additive 17.778 0.0185 3.8291 0.0472 37.16 0.0543 0.0978
hausdorff box_count_recall thresholded aps multiplicative 17.778 0.0185 1.8169 0.0396 37.1598 0.0543 0.0911
hausdorff box_count_recall pixelwise lac additive 17.778 0.0185 1.1577 0.051 38.134 0.0515 0.0946
hausdorff box_count_recall pixelwise lac multiplicative 17.778 0.0185 1.114 0.0506 38.134 0.0515 0.0953
hausdorff box_count_recall pixelwise aps additive 17.778 0.0185 1.1577 0.051 37.16 0.0543 0.0976
hausdorff box_count_recall pixelwise aps multiplicative 17.778 0.0185 1.114 0.0506 37.1598 0.0543 0.0982
hausdorff box_count_recall boxwise lac additive 17.778 0.0185 2.9928 0.0522 38.134 0.0515 0.0947
hausdorff box_count_recall boxwise lac multiplicative 17.778 0.0185 1.6957 0.0492 38.134 0.0515 0.0938
hausdorff box_count_recall boxwise aps additive 17.778 0.0185 2.9928 0.0522 37.16 0.0543 0.0978
hausdorff box_count_recall boxwise aps multiplicative 17.778 0.0185 1.6957 0.0492 37.1599 0.0543 0.0967
lac box_count_threshold thresholded lac additive 25.5884 0.022 21.6022 0.0408 0.9939 0.0508 0.0874
lac box_count_threshold thresholded lac multiplicative 25.5884 0.022 18.7336 0.0488 0.9939 0.0508 0.0942
lac box_count_threshold thresholded aps additive 25.5884 0.022 21.6022 0.0408 1.0066 0.0505 0.0871
lac box_count_threshold thresholded aps multiplicative 25.5884 0.022 18.7336 0.0488 1.0066 0.0505 0.0939
lac box_count_threshold pixelwise lac additive 25.5884 0.022 17.1966 0.0459 0.9939 0.0508 0.0855
lac box_count_threshold pixelwise lac multiplicative 25.5884 0.022 14.147 0.0501 0.9939 0.0508 0.0892
lac box_count_threshold pixelwise aps additive 25.5884 0.022 17.1966 0.0459 1.0066 0.0505 0.0852
lac box_count_threshold pixelwise aps multiplicative 25.5884 0.022 14.147 0.0501 1.0066 0.0505 0.0889
lac box_count_threshold boxwise lac additive 25.5884 0.022 20.4218 0.0431 0.9939 0.0508 0.082
lac box_count_threshold boxwise lac multiplicative 25.5884 0.022 16.2098 0.05 0.9939 0.0508 0.0878
lac box_count_threshold boxwise aps additive 25.5884 0.022 20.4218 0.0431 1.0066 0.0505 0.0816
lac box_count_threshold boxwise aps multiplicative 25.5884 0.022 16.2098 0.05 1.0066 0.0505 0.0876
lac box_count_recall thresholded lac additive 17.778 0.0185 20.0357 0.0352 1.2434 0.0482 0.0806
lac box_count_recall thresholded lac multiplicative 17.778 0.0185 20.0441 0.0404 1.2434 0.0482 0.0855
lac box_count_recall thresholded aps additive 17.778 0.0185 20.0357 0.0352 1.3256 0.0474 0.0797
lac box_count_recall thresholded aps multiplicative 17.778 0.0185 20.0441 0.0404 1.3256 0.0474 0.0846
lac box_count_recall pixelwise lac additive 17.778 0.0185 15.8575 0.0427 1.2434 0.0482 0.0826
lac box_count_recall pixelwise lac multiplicative 17.778 0.0185 15.1964 0.0457 1.2434 0.0482 0.0852
lac box_count_recall pixelwise aps additive 17.778 0.0185 15.8575 0.0427 1.3256 0.0474 0.082
lac box_count_recall pixelwise aps multiplicative 17.778 0.0185 15.1964 0.0457 1.3256 0.0474 0.0849
lac box_count_recall boxwise lac additive 17.778 0.0185 18.7786 0.0395 1.2434 0.0482 0.0785
lac box_count_recall boxwise lac multiplicative 17.778 0.0185 17.596 0.044 1.2434 0.0482 0.0822
lac box_count_recall boxwise aps additive 17.778 0.0185 18.7786 0.0395 1.3256 0.0474 0.078
lac box_count_recall boxwise aps multiplicative 17.778 0.0185 17.596 0.044 1.3256 0.0474 0.0819
giou box_count_threshold thresholded lac additive 25.5884 0.022 35.5329 0.0224 44.5444 0.0482 0.0697
giou box_count_threshold thresholded aps additive 25.5884 0.022 35.5329 0.0224 43.6363 0.0491 0.0707
giou box_count_threshold pixelwise lac additive 25.5884 0.022 31.4165 0.0359 44.5444 0.0482 0.0815
giou box_count_threshold pixelwise lac multiplicative 25.5884 0.022 158.2964 0.0263 44.5444 0.0482 0.0738
giou box_count_threshold pixelwise aps additive 25.5884 0.022 31.4165 0.0359 43.6364 0.0491 0.0822
giou box_count_threshold pixelwise aps multiplicative 25.5884 0.022 158.2964 0.0263 43.6363 0.0491 0.0748
giou box_count_threshold boxwise lac additive 25.5884 0.022 35.0271 0.0305 44.5444 0.0482 0.0764
giou box_count_threshold boxwise lac multiplicative 25.5884 0.022 183.4134 0.0239 44.5444 0.0482 0.0715
giou box_count_threshold boxwise aps additive 25.5884 0.022 35.0271 0.0305 43.6364 0.0491 0.0772
giou box_count_threshold boxwise aps multiplicative 25.5884 0.022 183.4134 0.0239 43.6363 0.0491 0.0724
giou box_count_recall thresholded lac additive 17.778 0.0185 31.8864 0.0144 44.6575 0.0458 0.0596
giou box_count_recall thresholded aps additive 17.778 0.0185 31.8864 0.0144 44.4713 0.0456 0.0593
giou box_count_recall pixelwise lac additive 17.778 0.0185 28.2407 0.0279 44.6575 0.0458 0.0716
giou box_count_recall pixelwise lac multiplicative 17.778 0.0185 165.6919 0.0114 44.6575 0.0458 0.0566
giou box_count_recall pixelwise aps additive 17.778 0.0185 28.2407 0.0279 44.4707 0.0456 0.0712
giou box_count_recall pixelwise aps multiplicative 17.778 0.0185 165.6919 0.0114 44.4713 0.0456 0.0565
giou box_count_recall boxwise lac additive 17.778 0.0185 31.4174 0.0204 44.6575 0.0458 0.0646
giou box_count_recall boxwise lac multiplicative 17.778 0.0185 192.0147 0.01 44.6575 0.0458 0.0553
giou box_count_recall boxwise aps additive 17.778 0.0185 31.4174 0.0204 44.4707 0.0456 0.0642
giou box_count_recall boxwise aps multiplicative 17.778 0.0185 192.0147 0.01 44.4713 0.0456 0.0551
TABLE VI: Experimental Results using the DETR-101 model for $\alpha^{\mathrm{tot}}=0.2$